# TC-Light: Temporally Consistent Relighting for Dynamic Long Videos

### v1

@@ Abstract
Editing illumination in long videos with complex dynamics has significant value in various downstream tasks, including visual content creation and manipulation, as well as data scaling up for embodied AI through sim2real and real2real transfer. Nevertheless, existing video relighting techniques are predominantly limited to portrait videos or fall into the bottleneck of temporal consistency and computation efficiency. In this paper, we propose TC-Light, a novel paradigm characterized by the proposed two-stage post optimization mechanism. Starting from the video preliminarily relighted by an inflated video relighting model, it optimizes appearance embedding in the first stage to align global illumination. Then it optimizes the proposed canonical video representation, i.e., Unique Video Tensor (UVT), to align fine-grained texture and lighting in the second stage. To comprehensively evaluate performance, we also establish a long and highly dynamic video benchmark. Extensive experiments show that our method enables physically plausible relighting results with superior temporal coherence and low computation cost. The code and video demos are available at https://dekuliutesla.github.io/tclight/.
@@ 1 Introduction
Lighting and its interaction with both real and synthetic environments fundamentally shapes how humans—and embodied agents—perceive the world. The ability to relight camera-captured image sequences, especially in complex, highly dynamic scenes, is invaluable for a range of downstream applications, spanning from filmmaking Richardt_Stoll_Dodgson_Seidel_Theobalt_2012 to augmented reality li2022physically . Crucially, this capability also benefits embodied AI: by adapting the illumination distribution in video while preserving intrinsic image details, it helps bridge the sim-to-real gap and enable real-to-real transfer, thus unlocking access to massive high-quality data that is essential for scaling up embodied intelligence.
Despite its importance, video relighting remains a highly challenging problem, particularly when camera motion is highly dynamic and foreground objects frequently enter and exit scenes, as shown in Fig. 1. Most existing relighting techniques zhang2025scaling ; li2022physically ; jin2024neural_gaffer ; ren2024relightful ; kim2024switchlight ; wang2023sunstage are tailored for static images. As shown in Sec. 4.2, naively inflating them to a video model with existing zero-shot strategies struggles to balance the consistency and quality. Moreover, the considerable training cost and scarcity of video lighting datasets hinder fine-tuning a pretrained model for this task. Besides, though video relighting models are emerging, they are either restricted on portrait video zhang2021neural ; choi2024personalized ; cai2024real , or burdened by considerable computation overhead zhou2025light ; fang2025relightvid on long video, as validated in Sec. 4.2.

### v2

# TC-Light: Temporally Coherent Generative Rendering for Realistic World Transfer

@@ Abstract
Illumination and texture editing are critical dimensions for world-to-world transfer, which is valuable for applications including sim2real and real2real visual data scaling up for embodied AI. Existing techniques generatively re-render the input video to realize the transfer, such as video relighting models and conditioned world generation models. Nevertheless, these models are predominantly limited to the domain of training data (e.g., portrait) or fall into the bottleneck of temporal consistency and computation efficiency, especially when the input video involves complex dynamics and long durations. In this paper, we propose TC-Light, a novel generative renderer to overcome these problems. Starting from the video preliminarily relighted by an inflated video relighting model, it optimizes appearance embedding in the first stage to align global illumination. Then it optimizes the proposed canonical video representation, i.e., Unique Video Tensor (UVT), to align fine-grained texture and lighting in the second stage. To comprehensively evaluate performance, we also establish a long and highly dynamic video benchmark. Extensive experiments show that our method enables physically plausible re-rendering results with superior temporal coherence and low computation cost. The code and video demos are available at https://dekuliutesla.github.io/tclight/.
@@ 1 Introduction
Lighting and its interaction with both real and synthetic environments fundamentally shapes how humans—and embodied agents—perceive the world. The ability to re-render the illumination and texture (or so-called relighting) of camera-captured image sequences, especially in complex, highly dynamic scenes, is critically valuable for various world-to-world transfer use cases. For instance, by re-rendering the CG-simulated or realistic video data used to train embodied agents, it can bridge the sim-to-real gap and enable real-to-real transfer, thus unlocking access to massive high-quality data that is essential for stepping towards embodied intelligence.
Despite its importance, the video illumination re-rendering remains a highly challenging problem, particularly when camera motion is highly dynamic and foreground objects frequently enter and exit scenes, which are common cases for embodied agents, as shown in Fig. 1. Most existing generative relighting techniques zhang2025scaling ; li2022physically ; jin2024neural_gaffer ; ren2024relightful ; kim2024switchlight ; wang2023sunstage are tailored for static images. As shown in Sec. 4.2, naively inflating them to a video model with existing zero-shot strategies struggles to balance the consistency and quality. Moreover, the considerable training cost and scarcity of video lighting datasets hinder fine-tuning a pretrained model for this task. Besides, though generative video relighting and world generation models are emerging, they are either restricted on domain of training data zhang2021neural ; choi2024personalized ; cai2024real ; alhaija2025cosmos or burdened by considerable computation overhead zhou2025light ; fang2025relightvid on long video, as validated in Sec. 4.2.